OpenStudy (yacoub1993):

In a 1-mile race, the winner crosses the finish line 10 feet ahead of the second-place runner and 23 feet ahead of the third-place runner. Assuming that each runner maintains a constant speed throughout the race, by how many feet does the second-place runner beat the third-place runner? (5280 feet in 1 mile.) A. -13.06 ft B. 13.02 ft C. 3.01 ft D. -10.04 ft

OpenStudy (yacoub1993):

@Directrix

Directrix (directrix):

Let's think.

OpenStudy (yacoub1993):

how do we start @Directrix

Directrix (directrix):

I am writing two equations and made an error I am correcting. These others are writing 23 - 10 as the answer but I am not yet convinced of that.

OpenStudy (yacoub1993):

ok take your time

Directrix (directrix):

F = 23 + T where F is the first place winner and T is the third place winner F = 10 + S where S is the second place winner.

Directrix (directrix):

So, S = F - 10 T = F - 23 ----------- Subtracting to get S - T (distance from second to third), the two Fs add out, leaving this: -10 - (-23) = ? @Yacoub1993 I guess it should have been obvious that those others were correct. :)

OpenStudy (yacoub1993):

A. -13.06 ft B. 13.02 ft C. 3.01 ft D. -10.04 ft

Directrix (directrix):

@Yacoub1993 What did you get for the answer option?

OpenStudy (yacoub1993):

there is some minor mistakes you did i think

Directrix (directrix):

I didn't see a minor mistake. I think the answer options are off. But, there is only one answer option which could be anywhere near correct. Do you see it?

OpenStudy (yacoub1993):

yes 13.02

Directrix (directrix):

It is the answer to this: -10 - (-23) =

OpenStudy (yacoub1993):

yes 13.02 is the nearest answer

Directrix (directrix):

I agree with that, the B option.